OpenStudy (anonymous):

what is the equation of the directrix of a parabola defined by the equation y=x^2-2x-3

OpenStudy (jdoe0001):

hmm have you covered parabolas yet?

OpenStudy (anonymous):

Yes but I don't fully understand it

OpenStudy (jdoe0001):

hmm ok.... are you supposed to get the vertex.... by factoring, or completing the square or graphing?

OpenStudy (jdoe0001):

well... you're needing the directrix.... thus why I mentioned the vertex since once you find the vertex form of the parabola, the directrix is also found

OpenStudy (anonymous):

Ok do I have to factor it out

OpenStudy (jdoe0001):

do yo know what a perfect square, or "perfect square trinomial" is?

OpenStudy (jdoe0001):

nope, you don't, but you could

OpenStudy (anonymous):

I heard of perfect square

OpenStudy (jdoe0001):

ok.... so \(\large \begin{array}{cccccllllll} {\color{brown}{ a}}^2& + &2{\color{brown}{ a}}{\color{blue}{ b}}&+&{\color{blue}{ b}}^2\\ \downarrow && &&\downarrow \\ {\color{brown}{ a}}&& &&{\color{blue}{ b}}\\ &\to &({\color{brown}{ a}} + {\color{blue}{ b}})^2&\leftarrow \end{array}\qquad % perfect square trinomial, negative middle term \begin{array}{cccccllllll} {\color{brown}{ a}}^2& - &2{\color{brown}{ a}}{\color{blue}{ b}}&+&{\color{blue}{ b}}^2\\ \downarrow && &&\downarrow \\ {\color{brown}{ a}}&& &&{\color{blue}{ b}}\\ &\to &({\color{brown}{ a}} - {\color{blue}{ b}})^2&\leftarrow \end{array}\) does that ring a bell?

OpenStudy (anonymous):

Yes looks familiar

OpenStudy (jdoe0001):

ok so let us "complete the square so let's see, we have \(\bf y=x^2-2x-3\implies y+3=(x^2-2x+{\color{red}{ \square }}^2)\) so...what do you think we're missing there, to gert a "perfect square trinomial" ?

OpenStudy (anonymous):

Umm is it 2

OpenStudy (jdoe0001):

2 ok well, notice, the middle term of the perfect square trinomial is 2 * the other two terms on its sides, without the exponent so 2 * x * 2 = 2x? ^ middle term so... guess not..... thus is not 2 t hen

OpenStudy (anonymous):

Is the missing term suppose to equal 3

OpenStudy (jdoe0001):

notice the middle term is " 2x " that means that \(\bf 2\cdot x\cdot {\color{red}{ \square }}=2x\) if we use 3 for the missing one 2 * x * 3 \(\ne 2x\)

OpenStudy (anonymous):

So its 1

OpenStudy (jdoe0001):

well... let's see 2 * x * 1 = 2x \(\Large \checkmark\) so now we know is 1 so keep in mind that all we're doing is borrowing from our good fellow Mr Zero, 0 thus, if we add \(1^2\) we also have to subtract \(1^2\) thus

OpenStudy (jdoe0001):

\(\bf y=x^2-2x-3\implies y+3=(x^2-2x+{\color{red}{ 1 }}^2)-{\color{red}{ 1 }}^2 \\ \quad \\ y+3(x-1)^2-1\implies y+4=(x-1)^2\implies 1(y+{\color{blue}{ (-4)}})=(x-{\color{brown}{ 1}})^2 \\ \quad \\ \begin{array}{llll} 4{\color{purple}{ p}}(y-{\color{blue}{ k}})=(x-{\color{brown}{ h}})^2\\ \end{array} \qquad \begin{array}{llll} vertex\ ({\color{brown}{ h}},{\color{blue}{ k}})\\ {\color{purple}{ p}}=\textit{distance from vertex to }\\ \qquad \textit{ focus or directrix} \end{array}\)

OpenStudy (anonymous):

After I plug in the values how do I solve it

OpenStudy (jdoe0001):

so our vertex is at ( 1, -4) the template for the vertex form, uses ( y -h) thus (y+4) is really ( y - (-4)) = (y+4) and notice the coefficient in front of it , is 1 so we know that "4p" is equals to 1 4p =1 p = 1/4 so the vertex is at (1, -4) and the directrix is 1/4 units from there the parabola has a leading term coefficient that is positive, that means the parabola is going upwards so that means the directrix is below the vertex